# Mixtera: A Data Plane for Foundation Model Training

### v2

@@ 5.2.2. Chunk generation
Chunk distribution. In distributed training, it is important to guarantee that all nodes within the same data parallel group operate on exactly the same input tensors. Mixtera’s ChunkDistributor wraps around the chunk generation component, and hands out chunks correctly to the training nodes, i.e., the same chunks in the same order to nodes within the same group, and different chunks to nodes in different groups for data parallelism. To this end, the clients need to register at the server with their respective node and group identifiers. To avoid redundant serialization overhead, the distributor caches chunks in serialized form until all nodes in a group have received them.
@@ 5.2.3. Networking
We implement a TCP-based client-server protocol. The server uses Python’s asyncio framework to handle multiple concurrent client connections. The protocol is message-based, with each message consisting of a task identifier followed by task-specific payload data. Tasks for example include the execution of a query or sending out a new chunk to a client. To handle network issues gracefully, the client implementation includes automatic reconnection with exponential backoff and configurable timeouts. Larger data transfers, such as chunk transmission, use efficient streaming to avoid memory issues.
@@ 5.3. Client-Side Reading
Mixtera’s client-side abstractions offer a generator that yields data samples for a query at the server. This generator internally follows a two-level nested iteration pattern: an outer iteration over chunks and an inner iteration over samples within each chunk. The outer iteration hides the complexity of network transfer and chunk generation, while the inner iteration hides the complexity of going from pointers in the chunk to actual samples. We discuss this inner step, i.e., how, given a chunk, we yield sample payloads.
@@
Tokenized mixture mode. This mode extends the strict window mode. It wraps the active iterators with a tokenizing iterator that takes the incoming string samples, tokenizes them, and yields tokenized samples (integer lists) with the correct sequence length. By setting the window size equal to the chunk size we guarantee that each chunk at least yields one window of tokenized samples. This mode is important for handling imbalanced collection with significantly varying text lengths. For example, in the The Pile (Gao et al., 2020), an average sample from the Books3 domain is much longer than one from the PubMed Abstracts domain.
Without tokenization, one string sample may correspond to potentially order-of-magnitudes more tokenized samples, which can heavily disturb the actual mixture seen during training. While this mode ensures precise mixture ratios at the token level, it may result in partial utilization of longer samples. This is an inherent issue of unbalanced datasets, and while Mixtera provides flexibility to handle it, the best approach is to process the datasets such that samples are (roughly) of similar size.
@@
Dataset abstractions. Training frameworks typically require specific dataset interfaces that supports multiprocessing with worker processes. Besides a general purpose interface, Mixtera offers a class extending torch’s IterableDataset, and a class compatible with the huggingface API. Each worker process at each node operates on its own chunk.
@@ 5.4. Framework Integration
Checkpointing. Mixtera’s API offers a function to be called on checkpoint. In order to restore from checkpoint, we need to know (i) which chunks have been handed out to which nodes and (ii) which data loader worker processes have yielded how many samples for each node. Mixtera implements (i) using the ChunkDistributor, which caches the query and the current state on checkpoint and is able to restore the in-memory state of the iterators for chunk generation based on this information. For (ii) the MixteraTorchDataset uses a shared memory segment to share with the main training process the status of the data loader workers. When we restore a checkpoint, we restore the state at the server, hand out the last chunks to each worker that they were working on, and then at the workers discard the previously yielded samples.
@@ 6. Evaluation
We explore the first question in a dynamic mixture case study on LLMs (Section 6.1) and a static mixture case study for VLMs (Section 6.2); the second question is explored in throughput benchmarks (Section 6.3). We conduct our experiments on the Clariden partition of the Alps supercomputer of the Swiss National Supercomputer Centre. Clariden consists of HPE Cray Supercomputing EX254n blades, each hosting two Quad GH200 nodes. Each node contains 4 interconnected groups of a Grace CPU with 72 cores, 128 GB DRAM, and a H100 Hopper GPU with 96 GB of HBM. The nodes are connected using 200 Gb/s HPE Slingshot interconnect. We refer to Fusco et al. (Fusco et al., 2024) for a more detailed analysis of the Alps supercomputer. The machines run Ubuntu Server 24.01 LTS with kernel 5.14.21, and we build upon the NVIDIA NGC 25.01 container with Python 3.12, a nightly build of PyTorch 2.6, NVIDIA driver 550.54.15 and CUDA 12.8. We add support for Mixtera and other data loaders on our fork of torchtitan (commit d989842)222Available at https://github.com/eth-easl/torchtitan-mixtera. (Liang et al., 2024). Torchtitan is part of the Pytorch ecosystem and straightforward to set up. We use Llama3-like models whose configurations can be found in Table 2. The 162M and 1.6B models are based on Jiang et al. (Jiang et al., 2024), while 3.6B follows Meta’s Llama 3.2 model. They do not have the same parameter count as torchtitan does not tie the weight embeddings. Our training and benchmarking data is based on The Pile (Gao et al., 2020). We heuristically split long samples with more than 3000 words, with max. 10 k samples per file.
@@ 6.1. Dynamic Mixing using ADO
We first demonstrate how to implement dynamic mixing algorithms in Mixtera, taking the ADO algorithm (Section 2.1.1) as an example. We show how dynamic mixing can improve model performance during pre-training. Dynamic mixing algorithms are an active area of research (Jiang et al., 2024; Chen et al., 2024a, 2023; Li et al., 2025; Albalak et al., 2023) and Mixtera facilitates exploration.
@@
Model Steps Mixture HellaSwag WinoGrande ARC-E ARC-C LogiQA2 Lambada (OAI) OpenBookQA PIQA SlimP. PPL Pile PPL 162M 15 k ADO 0.28 0.49 0.43 0.20 0.23 0.29 0.19 0.61 58.19 66.88 Default 0.28 0.52 0.44 0.19 0.23 0.29 0.18 0.61 59.13 69.52 Natural 0.28 0.52 0.41 0.18 0.23 0.26 0.16 0.61 57.54 69.85 30 k ADO 0.29 0.50 0.44 0.19 0.23 0.31 0.17 0.62 52.62 61.82 Default 0.28 0.52 0.44 0.20 0.23 0.34 0.16 0.62 53.55 64.39 Natural 0.29 0.54 0.44 0.19 0.24 0.31 0.16 0.63 52.78 65.03 1B 15 k ADO 0.37 0.54 0.51 0.23 0.25 0.52 0.20 0.70 30.07 32.20 Default 0.35 0.54 0.52 0.23 0.22 0.47 0.17 0.68 29.34 33.02 Natural 0.35 0.54 0.52 0.23 0.23 0.50 0.17 0.68 31.31 34.40 30 k ADO 0.40 0.55 0.55 0.26 0.23 0.54 0.23 0.71 25.89 29.05 Default 0.38 0.56 0.56 0.24 0.22 0.56 0.21 0.70 25.10 28.65 Natural 0.38 0.56 0.55 0.25 0.22 0.57 0.21 0.69 26.55 29.83 3B 15 k ADO 0.42 0.56 0.57 0.25 0.24 0.59 0.21 0.71 25.78 23.00 Default 0.38 0.54 0.52 0.25 0.24 0.50 0.20 0.68 26.21 24.73 30 k ADO 0.45 0.59 0.60 0.28 0.24 0.59 0.24 0.73 22.22 20.52 Default 0.42 0.56 0.59 0.26 0.24 0.57 0.23 0.70 22.52 21.79
Mixtera implementation. The original ADO implementation updates the current mixture after every step and samples the next batch based on this distribution. The distribution is updated based on the per-domain loss from the previous step. This does not fully match Mixtera, which can only use a new mixture when generating a new chunk. Each chunk then may yield several batches of data with the same mixture. We still send the per-domain losses on every training step at the client to update ADO’s internal state at the server. Whenever a new chunk is generated, the current mixture from ADO is queried (Algorithm 1), and the server generates a chunk according to that mixture. This introduces some slack since we do not use a new mixture on every batch. However, due to the stochastic nature of sampling, we do not find this to be an issue.
The only change needed in training clients is the implementation of a per-domain loss. For this, the loss function (e.g., cross-entropy) needs to be called without reduction, which gives a loss per token. As Mixtera provides which token belongs to which domain, we can aggregate the losses per domain. We then perform an all-reduce operation across all training nodes to get the global per-domain losses and send this to the server.
While the original implementation of ADO is fully tied to their training framework and data setup, Mixtera is agnostic to the training framework. While developing, we switched from nanotron (HuggingFace, 2025) to torchtitan (Liang et al., 2024). No changes in Mixtera were required. Only the per-domain loss module at the training framework is required. This showcases the benefit of having a system like Mixtera that decouples the mixing from the training framework.
@@
Training setup. We base our setup on Jiang et al. (Jiang et al., 2024). In addition to their 162M and 1.6B architectures, to test how ADO scales to the latest model architecture, we try a 3.6B model (Table 2) following Llama-3.2-3B from Meta, including its tokenizer. The 162M and 1.6B models use the EleutherAI/GPT-NeoX-20B tokenizer. All models use a sequence length of 2048. For ADO, we follow the codebase and discard the first 500 steps for fitting the scaling laws, start with fitting the scaling laws at step 1 000, and then re-fit the laws every 1 000 steps with a subsampling frequency of 10. We also follow the codebase and “use the same step size” for all domains. Instead of using the count of how often a domain has been sampled to fit the scaling laws like in the paper, we follow the ADO codebase and average the total sample counts evenly across all domains. We train all models with non-strict token-level mixtures, a learning rate of 0.001 with a linear warmup for 500 steps and linear decay until the final step, and the AdamW optimizer, for 30 000 steps. For 162M/1.6B, we use 64 GPUs with a microbatch size of 32, and for 3.6B we use 128 GPUs with a microbatch size of 16, resulting in a global batch size of 2048 and total 125 B tokens. As mixtures, we test ADO, the default weights as in the DoReMi paper (Xie et al., 2024), and on 162M/1.6B, the “natural” weights by Jiang et al. (Jiang et al., 2024) which are optimized for the NeoX tokenizer.
@@
Evaluation metrics. We follow Jiang et al. (Jiang et al., 2024) and report both downstream task performance as well as perplexity. For downstream tasks, we report performance on HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), ARC-Easy and ARC-Challenge (Clark et al., 2018), LogiQA2 (Liu et al., 2023), Lambada OpenAI (Paperno et al., 2016), OpenBookQA (Mihaylov et al., 2018), and PIQA (Bisk et al., 2020). For perplexity, we report the average unweighted token perplexity on (i) the validation set of The Pile (Gao et al., 2020), and on (ii) SlimPajama (Soboleva et al., 2023) as a dataset we did not train on. We collect all metrics using EleutherAI’s lm-eval-harness (Gao et al., 2023), and use the unnormalized accuracy.
ADO algorithm performance overview. Table 3 shows the performance of all models and mixtures for a checkpoint after 15 k and 30 k steps. We mark in bold the best value within a model/step group. In contrast to Jiang et al. (Jiang et al., 2024), we find that on the 162M model, the default mixture is quite competitive across benchmarks. For the final checkpoint (30 k), ADO achieves the best (lowest) perplexity on both SlimPajama and The Pile, but it does not consistently beat the static mixtures as reported in the ADO paper. However, analysis of the textual outputs reveals that the quality of generated text from this small model is limited across all configurations.
For the larger 1B model, ADO is particulary strong on the intermediate checkpoint, which shows an improvement in time-to-accuracy. For the final checkpoint, on some benchmarks, the static mixtures perform on-par or marginally better than ADO, but on others, such as HellaSwag, ADO clearly beats the static mixtures.
@@
On the 3B model, ADO beats the static default mixture on both checkpoints across all benchmarks. While Jiang et al. (Jiang et al., 2024) report that ADO performs best on the smaller scale model, we find that ADO performs better on larger models. Since ADO’s official repository is tightly coupled with the training framework, even after corresponding with the authors, we were not able to identify the root cause of this, partly also because their code is bound to training on specific cloud instances. This shows that a common codebase for data mixing decoupled from the training framework is beneficial and help developers integrate the latest algorithms into their setups (Section 3).
TODO!
@@
Performance over time. To demonstrate how different tasks behave over the training, we show the performance of the 1B model on HellaSwag, OpenBookQA, and ARC-Easy for all training checkpoints in Figure 5. Every benchmark exhibits different behavior. For HellaSwag, we can clearly see that ADO consistently increases its margin over the static mixtures. For OpenBookQA, ADO performs well especially in the intermediate checkpoints, and the static mixtures greatly improve towards the end. For ARC-Easy, all mixtures behave similarly. This motivates future research on data mixing using Mixtera. For example, we might be able to use intermediate evaluations instead of loss to dynamically adjust the mixture.
TODO!
@@
Mixture over time. We showcase the mixture over time for the six largest domains in Figure 6. Around step 21 k, we run out of data for the Books3 domain, and Mixtera’s best-effort algorithm re-distributes the weight proportionally to the other domains. This happens because The Pile’s samples are very imbalanced in size, i.e., the average sample in Books3 has around 538 KiB, while the average Pile-CC sample has 4 KiB. While we split long samples heuristically, there still is some waste due to token-level mixtures. Unlike Jiang et al. (Jiang et al., 2024), we do not observe that other models/tokenizers lead to different mixtures. Notably, the more parameters, the higher the weight of Books3, and the lower the weight of PubMed Central. On the 3.6B model, OpenWebText2’s weight even surpasses PubMed Central’s weight before step 5 000.
@@
Takeaways. Dynamic mixture algorithms can improve model accuracy. ADO scales beyond the 1.6B model tested in the original paper, beating the default weights on the 3.6B model on all benchmarks. Our experiments also demonstrate that a synchronous algorithm, which uses a new mixture at each training step, adapts to Mixtera’s chunking system. We observe some discrepancies on the 162M model, which highlights the importance of having a common, open-source data mixing platform like Mixtera that facilities debugging and reproducibility.

### v3

@@ 5.2.2. Chunk generation
Chunk distribution. In distributed training, it is important to guarantee that all nodes within the same data parallel group operate on the same input tensors. Mixtera’s ChunkDistributor wraps around the chunk generation component, and hands out chunks correctly to the training nodes, i.e., the same chunks in the same order to nodes within the same group, and different chunks to nodes in different groups for data parallelism. To this end, the clients need to register at the server with their respective node and group identifiers. To avoid redundant serialization overhead, the distributor caches chunks in serialized form until all nodes in a group have received them.
@@ 5.2.3. Networking
We implement a TCP-based client-server protocol. The server uses Python’s asyncio framework to handle multiple concurrent client connections. The protocol is message-based, with each message consisting of a task identifier followed by task-specific payload data. Tasks, for example, include the execution of a query or sending out a new chunk to a client. To handle network issues gracefully, the client implementation includes automatic reconnection with exponential backoff and configurable timeouts. Note that only small objects such as chunks and not actual training data is transferred via Mixtera, to avoid training bottlenecks.
@@ 5.3. Client-Side Reading
Mixtera’s client-side abstractions provide a generator that, given a chunk from the server, yields the actual sample payloads . This generator follows a two-level nested iteration pattern: an outer iteration over chunks and an inner iteration over samples within each chunk. The outer iteration hides the complexity of network transfer and chunk generation, while the inner iteration hides the complexity of going from pointers in the chunk to actual samples. The previous section discussed the outer step, and we now discuss this inner step, i.e., how, given a chunk, we yield sample payloads.
Sample granularity. A critical aspect of Mixtera’s client-side reading is understanding the sample unit. As defined in Section˜4.1, a sample is the atomic unit of training data as supplied by the user. The sample granularity is inherent to the ingested dataset . Mixtera is a read-only layer and does not modify or redefine this granularity. In practice, sample granularity varies significantly, from books to partial sentences.
Varying granularities creates a challenge for mixture enforcement. If samples come in different lengths, enforcing mixture proportions at the sample level may not reflect the actual mixture at the token level seen during training. A single long sample from one domain may contribute orders of magnitude more tokens than a short sample from another domain. To address this, Mixtera provides three processing modes that offer different trade-offs between mixture guarantee granularity and sample utilization.
@@
Tokenized mixture mode. This mode addresses the sample granularity issue identified above by enforcing mixture ratios at the token level rather than the sample level. It wraps the active iterators with a tokenizing iterator that takes the incoming string samples, tokenizes them, and yields tokenized samples (integer lists) with the correct sequence length. By setting the window size equal to the chunk size we guarantee that each chunk yields at least one window of tokenized samples. This ensures that the mixture is respected in terms of actual training tokens, regardless of the varying lengths of the underlying samples .
While this mode ensures precise mixture ratios at the token level, it may result in partial utilization of longer samples. This is an inherent issue of unbalanced datasets, and while Mixtera provides flexibility to handle it, the best approach is to process the datasets such that samples are (roughly) of similar size.
@@
Dataset abstractions. Training frameworks typically require specific dataset interfaces that support multiprocessing with worker processes. Besides a general-purpose interface, Mixtera offers a class extending torch’s IterableDataset, and a class compatible with the Hugging Face API. Each worker process at each node operates on its own chunk.
@@ 5.4. Framework Integration
Checkpointing. Mixtera’s API offers a function to be called on checkpoint. To restore from a checkpoint, we need to know (i) which chunks have been handed out to which nodes and (ii) which data loader worker processes have yielded how many samples for each node. Mixtera implements (i) using the ChunkDistributor, which caches the query and the current state on checkpoint and can restore the in-memory state of the iterators for chunk generation based on this information. For (ii) the MixteraTorchDataset uses a shared memory segment to share with the main training process the status of the data loader workers. When we restore a checkpoint, we restore the state at the server, hand out the last chunks to each worker that they were working on, and then at the workers discard the previously yielded samples.
@@ 6. Evaluation
We explore the first question in a dynamic mixture case study on LLMs (Section˜6.1) and a static mixture case study for VLMs (Section˜6.2). We explore the second question with throughput benchmarks (Section˜6.3).
Setup. We run experiments on HPE Cray Supercomputing EX254n blades, each hosting two Quad GH200 nodes. Each node contains 4 interconnected groups of a Grace CPU with 72 cores, 128 GB DRAM, and a H100 Hopper GPU with 96 GB of HBM. The nodes are connected using a 200 Gb/s HPE Slingshot interconnect. The machines run Ubuntu Server 24.04 LTS with kernel 5.14.21. We build on the NVIDIA NGC 25.01 container with Python 3.12, a nightly build of PyTorch 2.7, NVIDIA driver 550.54.15, and CUDA 12.8. We add support for Mixtera and other data loaders on our fork of Torchtitan (commit ae4e402)222Available at https://github.com/eth-easl/torchtitan-mixtera. (Liang et al., 2024). Torchtitan is part of the PyTorch ecosystem and straightforward to set up. We use Llama3-like models with configurations in Table˜2. The 162M and 1.6B models are based on Jiang et al. (Jiang et al., 2024), while 3.6B follows Meta’s Llama 3.2 model. They do not have the same parameter count as torchtitan does not tie the weight embeddings. Our training and benchmarking data is based on The Pile (Gao et al., 2020), a frequently-used data collection used for exploring data mixing (Jiang et al., 2024; Xie et al., 2024; Albalak et al., 2023). We split long samples with more than 1 500 words, with max. 20 k samples per file.
@@ 6.1. Dynamic Mixing using ADO
We demonstrate how to implement dynamic mixing algorithms in Mixtera and their impact on model performance, taking the ADO algorithm (Section˜2.1.1) as an example.
@@
Model Mixture HellaSwag WinoGrande ARC-E ARC-C Lambada (OAI) OpenBookQA PIQA SlimP. Perp. Pile Perp. 1.6B ADO 0.411 0.577 0.577 0.262 0.603 0.236 0.707 25.20 27.79 Default 0.380 0.549 0.543 0.243 0.551 0.218 0.695 26.55 30.72 Natural 0.383 0.556 0.559 0.247 0.562 0.214 0.698 26.04 29.94 3.6B ADO 0.449 0.626 0.601 0.276 0.624 0.244 0.732 22.58 22.26 Default 0.415 0.597 0.579 0.255 0.592 0.210 0.706 22.09 22.02 Natural 0.419 0.586 0.593 0.276 0.598 0.206 0.723 21.99 22.53
@@
Training setup. We test the 1.6B model from Jiang et al. (Jiang et al., 2024) with the EleutherAI/GPT-NeoX-20B tokenizer and a 3.6B model (Table˜2) following Llama-3.2-3B from Meta, including its tokenizer. We omit the results from the 162M model for brevity. We use a sequence length of 2048. For ADO, we follow the codebase and discard the first 500 steps for fitting the scaling laws, start with fitting them at step 1 000, and then re-fit the laws every 1 000 steps with a subsampling frequency of 10. We also follow the codebase and “use the same step size” for all domains, i.e., instead of using the count of how often a domain has been sampled to fit the scaling laws like in the paper, we average the total sample counts evenly across all domains. We train using non-strict token-level mixtures for 30 000 steps, using a learning rate of 0.001 with a linear warmup for 500 steps and linear cooldown for 3 000 steps, and the AdamW optimizer. For 1.6B, we use 64 GPUs with a microbatch size of 32, and for 3.6B we use 128 GPUs with a microbatch size of 16, resulting in a global batch size of 2048 and total 125 B tokens. As mixtures, we test the default weights as in DoReMi (Xie et al., 2024), the natural weights as in Jiang et al. (Jiang et al., 2024), and ADO initialized with the natural weights.
@@
Evaluation metrics. We follow Jiang et al. (Jiang et al., 2024) and report both downstream task performance as well as perplexity. For downstream tasks, we report performance on HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), ARC-Easy and ARC-Challenge (Clark et al., 2018), Lambada OpenAI (Paperno et al., 2016), OpenBookQA (Mihaylov et al., 2018), and PIQA (Bisk et al., 2020). For perplexity, we report the average unweighted token perplexity on (i) the validation set of The Pile (Gao et al., 2020), and on (ii) SlimPajama (Soboleva et al., 2023) as a dataset we did not train on. We collect all metrics using EleutherAI’s lm-eval-harness (Gao et al., 2023), and use the unnormalized accuracy.
ADO algorithm performance overview. Table˜3 shows the performance of all models and mixtures for the final checkpoint after learning rate cooldown. We mark in bold the best value within a model/step group. Generally, ADO beats the static mixtures across all downstream benchmarks. On 1.6B, it also has the best (lowest) perplexity on both SlimPajama and The Pile, while on 3.6B, the static mixtures have slightly lower perplexity. This shows that ADO leads to better performance for downstream tasks across scales, and demonstrates the usage of a dynamic mixing algorithm in Mixtera.
@@
Jiang et al. (Jiang et al., 2024) report that on the 1.6B model, ADO sometimes performs worse than static mixtures, which we do not confirm. Since ADO’s official repository is tightly coupled with the training framework, even after corresponding with the authors, we were not able to identify the root cause of this, partly also because their code is bound to training on specific cloud instances. Mixtera decouples the training framework from the mixing algorithm, which helps developers port existing algorithms to their setups (Section˜3).
@@
Performance over time. To demonstrate how different tasks behave over the training, we show the performance of the 1.6B model on HellaSwag, OpenBookQA, and ARC-Easy for all training checkpoints in Figure˜5. Every benchmark exhibits different behavior. For HellaSwag, ADO consistently increases its margin over the static mixtures. For OpenBookQA, ADO performs similarly to the natural mix in the intermediate checkpoints, and benefits a lot during the learning rate cooldown. For ARC-Easy, the mixtures perform similarly, with ADO having a small edge. This motivates future research on data mixing using Mixtera–for example, we might be able to use intermediate evaluations instead of loss to dynamically adjust the mixture.
@@
Mixture over time. We showcase the mixture over time for the six largest domains in Figure˜6. We show the mixture obtained on the 1.6B model, but unlike Jiang et al. (Jiang et al., 2024), we do not observe that other models/tokenizers lead to largely different mixtures. In all cases, the weight of GitHub and ArXiv rapidly decreases. Notably, the more parameters, the higher the weight of Books3, and the lower the weight of PubMed Central. On the 3.6B model, OpenWebText2’s weight surpasses PubMed Central’s weight before step 5 000, while for the 1.6B model, they only slowly approach.
Mixtera implementation. We implement ADO in Mixtera in ca. 800 LOC. The class only handles the core algorithm, while Mixtera hides the complexity of the actual mixing from the algorithm implementation. We implement it from scratch as the original implementation of ADO is fully tied to their training framework and data setup. The original ADO implementation updates the current mixture after every step and samples the next batch based on this distribution. This is a small discrepancy to Mixtera, which can only use a new mixture when generating a new chunk. Each chunk may then yield several batches of data with the same mixture. We still send the per-domain losses on every training step at the client to update ADO’s internal state at the server. Whenever a new chunk is generated, the current mixture from ADO is queried (Algorithm˜1), and the server generates a chunk according to that mixture. As we find in the experiments, this slight slack does not harm performance, due to the stochastic nature of sampling.
In order to use ADO, at the training nodes, the only change needed is the implementation of a per-domain loss. For this, the loss function (e.g., cross-entropy) needs to be called without reduction, which gives a loss per token. As Mixtera provides which token belongs to which domain, we can aggregate the losses per domain. We then perform an all-reduce operation across all training nodes to get the global per-domain losses and send this to the server.
During development, we switched from nanotron (HuggingFace, 2025) to torchtitan (Liang et al., 2024). Notably, as Mixtera is agnostic to the training framework, no changes in Mixtera were required. This showcases the benefit of having a system like Mixtera that decouples the mixing from the training framework.
@@
Takeaways. Dynamic mixtures can improve model accuracy. ADO scales beyond the 1.6B model tested in the original paper, and beats the default weights on all benchmarks. Our experiments also demonstrate that a synchronous algorithm, which uses a new mixture at each training step, adapts to Mixtera’s chunking system.